# Differentially Private Synthetic Data via Foundation Model APIs 2: Text

### v1

@@ Appendix B Additional Experimental Details
  • Yelp: Yelp data is a public benchmark providing reviews on businesses, and we used the preprocessed Yelp from (Yue et al., 2023). The number of train/val/test samples and label information in Tb. 9.
  • OpenReview: For OpenReview ICLR2023 data, we crawl the meta-data for each review using the OpenReview Python library, [^8] https://github.com/openreview/openreview-py and concatenate the fields “summary_of_the_paper”, “strength_and_weaknesses” and “summary_of_the_review” as one sample in our dataset. We group the two attributes – review area and recommendation – together as a combination, and drop the training samples from combinations that contain fewer than 50 training samples. The number of samples after such preprocessing and label information is provided in Tb. 9. The number of samples for each class is provided in Tb. 10 and Tb. 11.
  • PubMed: we use PubMed with abstracts of medical papers [^9] https://www.ncbi.nlm.nih.gov/ crawled by Yu et al. (2023) from 2023/08/01 to 2023/08/07. The number of train/val/test samples are reported in Tb. 9.
@@ B.2.1 Model and Hyperparameters
We consider four LLMs as data generators in Aug-PE via API-access: GPT-2 (Radford et al., 2019), GPT-2-Medium, GPT-2-Large, and GPT-3.5 (“gpt-35-turbo” hosted on Microsft Azure [^10] https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/models(OpenAI, 2022). We provide the default hyper-parameter setup for GPT-3.5 in Tb. 12 and GPT-2 series models in Tb. 13.
@@
System
You are required to write an example of review based on the provided Business Category and Review Stars that fall within the range of 1.0-5.0.
Given the area and final decision of a research paper, you are required to provide an example of the review consisting of the following content: 1. briefly summarizing the paper in 3-5 sentences; 2. listing the strengths and weaknesses of the paper in details; 3. briefly summarizing the review in 3-5 sentences.
Please act as a sentence generator for the medical domain. Generated sentences should mimic the style of PubMed journal articles, using a variety of sentence structures.
User
Business Category: $\mathrm{label\_1}$} | Review Stars: $\mathrm{label\_2}$} with keyword $\mathrm{subcategory}$}
Area: $\mathrm{label\_1}$} | Recommendation: $\mathrm{label\_2}$}
Suppose that you are a $\mathrm{writer}$}. Please write an abstract for a medical research paper:
@@
Business Category: $\mathrm{label\_1}$} | Review Stars: $\mathrm{label\_2}$} with keyword $\mathrm{subcategory}$}
Suppose that you are a $\mathrm{writer}$}. Write a paper review based on Area: $\mathrm{label\_1}$} | Recommendation: $\mathrm{label\_2}$}
Using a variety of sentence structures, write an abstract for a medical research paper:
@@
Yelp
Based on “Business Category: $\mathrm{label\_1}$} | Review Stars: $\mathrm{label\_2}$}”, please rephrase the following sentences {in a $\mathrm{selected\_tone}$}:
@@
OpenReview
Based on “Area: $\mathrm{label\_1}$} | Recommendation: $\mathrm{label\_2}$}”, please rephrase the following sentences {in a $\mathrm{selected\_tone}$}:
@@
PubMed
Please rephrase the following sentences {in a $\mathrm{selected\_tone}$} as an abstract for medical research paper:
@@
System
You are a helpful, pattern-following assistant.
User
Based on the Business Category and Review Stars, you are required to fill in the blanks in the Input sentences. If there are no blanks, you are required to output the original Input sentences.
@@
System
You are an AI assistant that helps people find information
User
Based on the area and final recommendation of a research paper, you are required to fill in the blanks for the input sentences {in a $\mathrm{selected\_tone}$}. If there is no blanks, please output the original input sentences.
@@
System
You are an AI assistant that helps people find information.
User
You are required to fill in the blanks with more details for the input medical abstract {in a $\mathrm{selected\_tone}$}. If there is no blanks, please output the original medical abstract.
@@ B.2.2 API Prompt Designs
We provide the prompts of RANDOM_API for all datasets in Tb. 14 for GPT-3.5 and Tb. 15 for other LLMs.
@@
We provide the prompts of VARIATION_API for GPT-2-series models in Tb. 16 and for GPT-3.5 in Tb. 17, Tb. 18 and Tb. 19.
@@
For DP-FT-Downstream, we report the hyperparameters for OpenReview and Yelp in Tb. 21, and PubMed in Tb. 20. For a target $\epsilon$, a noise multiplier is set as the smallest value such that DP-SGD can run the target number of steps.
@@ Appendix C Additional Experimental Results
C.1 Convergence of Text Length Distribution
@@
C.2 Efficiency in Terms of GPU Hours
@@
In Tb. 22, we provide a detailed breakdown of the GPU hours shown in Fig. 6. We consider the process of generating DP synthetic data given a private dataset. DP-FT-Generator (Yue et al., 2023) requires two steps: (1) finetuning a pretrained data generator with DP-SGD, and (2) generating samples from the finetuned data generator, whereas Aug-PE requires only one step (Alg. 1). In Tb. 22, we list the GPU hours of each step of each method. For Yue et al. (2023), we use the hyper-parameters in their Table 8.
@@
C.3 Comparison Between Aug-PE and PE
We compare Aug-PE against PE when using GPT-3.5 as the generator on three datasets. The results in Tb. 23 show that Aug-PE is always better than PE on PubMed for GPT-3.5. Moreover, Aug-PE is better for OpenReview Rating classification task and Yelp Rating classification task. As Aug-PE supports PE as a special case by changing the hyperparameters of $L$ and $K$, the practitioner can adjust those hyperparameters for a specific downstream task and find the best settings to generate synthetic data.
@@ C.3 Comparison Between Aug-PE and PE
C.4 Ablation Study on Variation API Prompt Design
@@
The results in Tb. 24 show that fill-in-the-blanks prompt (with few-shot demonstrations) yields better results for GPT-3.5. For GPT-2, paraphrasing can be an effective strategy. Although fill-in-blanks leads to high accuracy on Yelp Category classification task, we find that the generated texts have many unfilled blanks “__” upon inspection.
C.5 Leveraging Open-source LLMs as Generator for Aug-PE
@@
C.6 Effect of Rank-based Sampling
@@
We compare our proposed rank-based sampling (Alg. 1) against probability-based random sampling in the original PE (Alg. 1) across GPT-2, GPT-2-Medium and GPT-2-Large on three datasets. The results in Tb. 25 indicate that our proposed rank-based sampling (Alg. 1) consistently outperforms probability-based random sampling in the original PE (Alg. 1), due to the elimination of sample redundancy inherent in random sampling, as rank-based sampling exclusively selects the top $N_{\mathrm{syn}}$ samples.
C.7 Embedding Distribution Distance Between Real and Synthetic data
@@
C.8 Downstream Task Utility Under Various Synthetic Data Size
C.8.1 Utility on Yelp
We report the full results of downstream accuracy on Yelp in Tb. 26. We find that (1) when using the same base model for a fair comparison, we see that under DP settings, Aug-PE demonstrates competitive (or even better) utility on downstream classification tasks compared to fine-tuning. The scores are also close to that of the downstream algorithms trained on the real data under DP directly, demonstrating the promise of DP synthetic text as a tool for DP machine learning. (2) For large models like GPT-2-Large and GPT-2-Medium, more synthetic samples (e.g., 100k) from Aug-PE can enhance downstream utility. However, for GPT-2, sometimes 10k synthetic samples can lead to better downstream utility than 100k samples, which might be due to the low-quality data generated from the small model that hurts the performance.
@@ C.8.1 Utility on Yelp
C.8.2 Utility on OpenReview
We report the downstream accuracy on OpenReview in Tb. 27. The key observations are: (1) Under DP when using the same GPT-2/GPT-2-Medium/GPT-2-Large as the base model, Aug-PE achieve similar classification accuracy and classification accuracy compared with DP-FT-Generator. This again demonstrates that Aug-PE is a promising alternative to DP fine-tuning. (2) More synthetic samples lead to better area classification accuracy for the three GPT-2-series models, indicating that Aug-PE scales well with the synthetic sample size. Note that both Aug-PE and DP-FT-Generator do not perform well on review rating classification tasks across different data sizes, which shows the inherent limitation of GPT-2-series models – they may struggle to generate academic texts with correct sentiments. (3) Aug-PE with GPT-3.5 achieves better utility than Aug-PE with GPT-2-Large on both tasks with or without DP. This suggests that Aug-PE benefits from larger and more powerful LLMs. We expect that as the capability of LLMs quickly evolves, Aug-PE can be even more promising in the future. (4) However, there is still a gap between the results of Aug-PE under non-DP setting $\epsilon=\infty$ and the results on the original data. This suggests that even in the non-DP setting, Aug-PE is still not able to recover the distribution of the real data. This gap is unavoidable in the DP setting. We hypothesize that better hyper-parameter tunings (e.g., the variation degree) could lower the gap. We leave a more careful investigation of this issue to future work.
C.8.3 Utility on PubMed
@@
We report the next-word prediction accuracy on OpenReview of downstream model $\text{BERT}_{\text{Mini}}$ in Tb. 28 and $\text{BERT}_{\text{Small}}$ in Tb. 29 We find that (1) under the same GPT-2-series model as generator, Aug-PE underperforms DP-FT-Generator on PubMed. This is expected because Aug-PE relies on the knowledge within LLMs to generate high-quality texts without domain-specific finetuning, while GPT-2-series models might have limited exposure to biomedical literature (Radford et al., 2019). (2) With powerful LLMs like GPT-3.5, Aug-PE can outperform DP-FT-Generator under DP. (3) Additionally, more synthetic samples lead to better downstream classification accuracy for the three GPT-2-series models on PubMed.
C.9 Aug-PE Convergence under One Private Sample

### v2

@@ 5 Conclusion
Acknowledgements
The authors thank Xiang Yue, Janardhan Kulkarni and anonymous reviewers for their valuable feedback and suggestions. BL gratefully acknowledges the support from the National Science Foundation under grants No. 1910100, No. 2046726, No. 2229876, and the Alfred P. Sloan Fellowship.
@@ Appendix B Additional Experimental Details
  • Yelp: Yelp data is a public benchmark providing reviews on businesses, and we used the preprocessed Yelp from (Yue et al., 2023). The number of train/val/test samples and label information in Tb. 10.
  • OpenReview: For OpenReview ICLR2023 data, we crawl the meta-data for each review using the OpenReview Python library, [^8] https://github.com/openreview/openreview-py and concatenate the fields “summary_of_the_paper”, “strength_and_weaknesses” and “summary_of_the_review” as one sample in our dataset. We group the two attributes – review area and recommendation – together as a combination, and drop the training samples from combinations that contain fewer than 50 training samples. The number of samples after such preprocessing and label information is provided in Tb. 10. The number of samples for each class is provided in Tb. 11 and Tb. 12.
  • PubMed: we use PubMed with abstracts of medical papers [^9] https://www.ncbi.nlm.nih.gov/ crawled by Yu et al. (2023) from 2023/08/01 to 2023/08/07. The number of train/val/test samples are reported in Tb. 10.
@@ B.2.1 Model and Hyperparameters
We consider four LLMs as data generators in Aug-PE via API-access: GPT-2 (Radford et al., 2019), GPT-2-Medium, GPT-2-Large, and GPT-3.5 (“gpt-35-turbo” hosted on Microsft Azure [^10] https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/models(OpenAI, 2022). We provide the default hyper-parameter setup for GPT-3.5 in Tb. 13 and GPT-2 series models in Tb. 14.
@@
System You are required to write an example of review based on the provided Business Category and Review Stars that fall within the range of 1.0-5.0. Given the area and final decision of a research paper, you are required to provide an example of the review consisting of the following content: 1. briefly summarizing the paper in 3-5 sentences; 2. listing the strengths and weaknesses of the paper in details; 3. briefly summarizing the review in 3-5 sentences. Please act as a sentence generator for the medical domain. Generated sentences should mimic the style of PubMed journal articles, using a variety of sentence structures.
User Business Category: $\mathrm{label\_1}$} | Review Stars: $\mathrm{label\_2}$} with keyword $\mathrm{subcategory}$} Area: $\mathrm{label\_1}$} | Recommendation: $\mathrm{label\_2}$} Suppose that you are a $\mathrm{writer}$}. Please write an abstract for a medical research paper:
@@
Business Category: $\mathrm{label\_1}$} | Review Stars: $\mathrm{label\_2}$} with keyword $\mathrm{subcategory}$} Suppose that you are a $\mathrm{writer}$}. Write a paper review based on Area: $\mathrm{label\_1}$} | Recommendation: $\mathrm{label\_2}$} Using a variety of sentence structures, write an abstract for a medical research paper:
@@
Yelp Based on “Business Category: $\mathrm{label\_1}$} | Review Stars: $\mathrm{label\_2}$}”, please rephrase the following sentences {in a $\mathrm{selected\_tone}$}:
@@
OpenReview Based on “Area: $\mathrm{label\_1}$} | Recommendation: $\mathrm{label\_2}$}”, please rephrase the following sentences {in a $\mathrm{selected\_tone}$}:
@@
PubMed Please rephrase the following sentences {in a $\mathrm{selected\_tone}$} as an abstract for medical research paper:
@@
System You are a helpful, pattern-following assistant.
User Based on the Business Category and Review Stars, you are required to fill in the blanks in the Input sentences. If there are no blanks, you are required to output the original Input sentences.
@@
System You are an AI assistant that helps people find information.
User Based on the area and final recommendation of a research paper, you are required to fill in the blanks for the input sentences {in a $\mathrm{selected\_tone}$}. If there is no blanks, please output the original input sentences.
@@
System Please act as a sentence generator for the medical domain. Generated sentences should mimic the style of PubMed journal articles, using a variety of sentence structures.
User You are required to fill in the blanks with more details for the input medical abstract {in a $\mathrm{selected\_tone}$}. If there is no blanks, please output the original medical abstract.
@@ B.2.2 API Prompt Designs
We provide the prompts of RANDOM_API for all datasets in Tb. 15 for GPT-3.5 and Tb. 16 for other LLMs.
@@
We provide the prompts of VARIATION_API for GPT-2-series models in Tb. 17 and for GPT-3.5 in Tb. 18, Tb. 19 and Tb. 20.
@@
For DP-FT-Downstream, we report the hyperparameters for OpenReview and Yelp in Tb. 22, and PubMed in Tb. 21. For a target $\epsilon$, a noise multiplier is set as the smallest value such that DP-SGD can run the target number of steps.
@@ Appendix C Additional Experimental Results
C.1 Robustness Against Membership Inference Attacks
In this work, we focus on DP, a type of widely accepted privacy guarantee with profound theoretic backup which provides an upper bound for empirical membership privacy attacks. To better understanding empirical risk, we perform state-of-the-art membership inference attacks (MIAs) (Shokri et al., 2017) in the text domain.
We perform MIAs against the finetuned downstream models (which are fine-tuned on synthetic data for Aug-PE/ DP-FT-Generator; on real private data for DP-FT-Downstream). We randomly sample 4000 PubMed real private samples as members and 4000 PubMed test samples as non-members for evaluation. We report AUC (Area Under the Curve) to evaluate the risks of MIAs. We consider three types of MIAs: (1) PPL thresholds perplexity to predict membership (Carlini et al., 2021). (2) REFER computes the ratio of the log perplexity of the tested model against a reference model (Carlini et al., 2021). (3) LIRA uses the ratio of likelihood instead of log-perplexity (Carlini et al., 2022). LiRA assumes the availability of high-quality data distributed similarly to the training set, which was thought to be impractical (Tramèr et al., 2022). Therefore, we follow (Mattern et al., 2023) to use the pre-trained model as a reference.
The results in Tb. 3 show that Aug-PE generally exhibits lower MIA AUC scores compared to both DP-FT-Generator and DP-FT-Downstream models. This indicates a higher robustness to empirical privacy attacks, potentially due to the synthetic nature of the data used for downstream model finetuning, which inherently reduces the risk of overfitting to real private data.
C.2 Convergence of Text Length Distribution
@@
C.3 Efficiency in Terms of GPU Hours
@@
In Tb. 23, we provide a detailed breakdown of the GPU hours shown in Fig. 6. We consider the process of generating DP synthetic data given a private dataset. DP-FT-Generator (Yue et al., 2023) requires two steps: (1) finetuning a pretrained data generator with DP-SGD, and (2) generating samples from the finetuned data generator, whereas Aug-PE requires only one step (Alg. 1). In Tb. 23, we list the GPU hours of each step of each method. For Yue et al. (2023), we use the hyper-parameters in their Table 8.
@@
C.4 Comparison Between Aug-PE and PE
We compare Aug-PE against PE when using GPT-3.5 as the generator on three datasets. The results in Tb. 24 show that Aug-PE is always better than PE on PubMed for GPT-3.5. Moreover, Aug-PE is better for OpenReview Rating classification task and Yelp Rating classification task. As Aug-PE supports PE as a special case by changing the hyperparameters of $L$ and $K$, the practitioner can adjust those hyperparameters for a specific downstream task and find the best settings to generate synthetic data.
@@ C.4 Comparison Between Aug-PE and PE
C.5 Ablation Study on Variation API Prompt Design
@@
The results in Tb. 25 show that fill-in-the-blanks prompt (with few-shot demonstrations) yields better results for GPT-3.5. For GPT-2, paraphrasing can be an effective strategy. Although fill-in-blanks leads to high accuracy on Yelp Category classification task, we find that the generated texts have many unfilled blanks “__” upon inspection.
C.6 Leveraging Open-source LLMs as Generator for Aug-PE
@@
C.7 Effect of Rank-based Sampling
@@
We compare our proposed rank-based sampling (Alg. 1) against probability-based random sampling in the original PE (Alg. 1) across GPT-2, GPT-2-Medium and GPT-2-Large on three datasets. The results in Tb. 26 indicate that our proposed rank-based sampling (Alg. 1) consistently outperforms probability-based random sampling in the original PE (Alg. 1), due to the elimination of sample redundancy inherent in random sampling, as rank-based sampling exclusively selects the top $N_{\mathrm{syn}}$ samples.
C.8 Effect of Iteration $T$ on DP Utility
Tb. 6 presents the results on the non-DP setting, serving as an ablation study to underscore the role of private data in Aug-PE. In the DP setting, given a fixed privacy budget, a larger $T$ requires more noise, which may compromise the utility of the DP histogram. On the other hand, a larger $T$ allows for more iterations of sample improvement. We want to study the joint effect of these two factors.
We conducted experiments comparing the utility of the algorithm at $t=1$ and $t=10$ on three datasets under $\epsilon=4,2,1$. The results in Tbs. 27, 28 and 29 show that $t=10$ consistently yields better utility than $t=1$ across all three privacy budget levels, underscoring the effectiveness of Aug-PE’s iterative improvement mechanism. This finding suggests that, despite the increased noise, the algorithm can robustly preserve useful statistical properties in the DP histogram and generate high-quality texts under $t=10$.
Table 27: Effect of Aug-PE iteration $t$ on the DP utility under $\epsilon=4$.
Yelp OpenReview PubMed
$t=1$ 64.7 73.5 36.5 42.1 29.9 32.3
$t=10$ 67.8 74.6 43.5 44.6 30.3 32.5
Table 28: Effect of Aug-PE iteration $t$ on the DP utility under $\epsilon=2$.
Yelp OpenReview PubMed
$t=1$ 63.9 73.6 37.2 42.0 29.9 32.3
$t=10$ 67.4 74.3 42.8 44.5 30.2 32.5
Table 29: Effect of Aug-PE iteration $t$ on the DP utility under $\epsilon=1$.
Yelp OpenReview PubMed
$t=1$ 63.8 73.1 37.4 41.7 30.0 32.3
$t=10$ 66.8 74.7 41.9 43.1 30.1 32.4
C.9 Embedding Distribution Distance Between Real and Synthetic data
@@
C.10 Downstream Task Utility Under Various Synthetic Data Size
C.10.1 Utility on Yelp
We report the full results of downstream accuracy on Yelp in Tb. 30. We find that (1) when using the same base model for a fair comparison, we see that under DP settings, Aug-PE demonstrates competitive (or even better) utility on downstream classification tasks compared to fine-tuning. The scores are also close to that of the downstream algorithms trained on the real data under DP directly, demonstrating the promise of DP synthetic text as a tool for DP machine learning. (2) For large models like GPT-2-Large and GPT-2-Medium, more synthetic samples (e.g., 100k) from Aug-PE can enhance downstream utility. However, for GPT-2, sometimes 10k synthetic samples can lead to better downstream utility than 100k samples, which might be due to the low-quality data generated from the small model that hurts the performance.
@@ C.10.1 Utility on Yelp
C.10.2 Utility on OpenReview
We report the downstream accuracy on OpenReview in Tb. 31. The key observations are: (1) Under DP when using the same GPT-2/GPT-2-Medium/GPT-2-Large as the base model, Aug-PE achieve similar classification accuracy and classification accuracy compared with DP-FT-Generator. This again demonstrates that Aug-PE is a promising alternative to DP fine-tuning. (2) More synthetic samples lead to better area classification accuracy for the three GPT-2-series models, indicating that Aug-PE scales well with the synthetic sample size. Note that both Aug-PE and DP-FT-Generator do not perform well on review rating classification tasks across different data sizes, which shows the inherent limitation of GPT-2-series models – they may struggle to generate academic texts with correct sentiments. (3) Aug-PE with GPT-3.5 achieves better utility than Aug-PE with GPT-2-Large on both tasks with or without DP. This suggests that Aug-PE benefits from larger and more powerful LLMs. We expect that as the capability of LLMs quickly evolves, Aug-PE can be even more promising in the future. (4) However, there is still a gap between the results of Aug-PE under non-DP setting $\epsilon=\infty$ and the results on the original data. This suggests that even in the non-DP setting, Aug-PE is still not able to recover the distribution of the real data. This gap is unavoidable in the DP setting. We hypothesize that better hyper-parameter tunings (e.g., the variation degree) could lower the gap. We leave a more careful investigation of this issue to future work.
@@ C.10.2 Utility on OpenReview
C.10.3 Utility on PubMed
@@
We report the next-word prediction accuracy on OpenReview of downstream model $\text{BERT}_{\text{Mini}}$ in Tb. 32 and $\text{BERT}_{\text{Small}}$ in Tb. 33 We find that (1) under the same GPT-2-series model as generator, Aug-PE underperforms DP-FT-Generator on PubMed. This is expected because Aug-PE relies on the knowledge within LLMs to generate high-quality texts without domain-specific finetuning, while GPT-2-series models might have limited exposure to biomedical literature (Radford et al., 2019). (2) With powerful LLMs like GPT-3.5, Aug-PE can outperform DP-FT-Generator under DP. (3) Additionally, more synthetic samples lead to better downstream classification accuracy for the three GPT-2-series models on PubMed.
C.11 Comparision to Text-to-Text Privatization Approaches
This is an active line of research on text-to-text privatization techniques for generating differentially private text. We do not directly compare these methods in our main paper due to the key distinctions in privacy definitions:
  1. 1.
    Different privacy definitions. Our method adopts the standard $(\epsilon,\delta)$-DP defined over neighboring datasets. This contrasts with
    1. (a)
      Word-level metric DP (Feyisetan et al., 2020; Carvalho et al., 2023): a specific metric for measuring word distance needs to be written in the privacy notation, and privacy guarantee is defined over neighboring words;
    2. (b)
      Local DP (Mattern et al., 2022b; Utpala et al., 2023): privacy guarantee is defined over neighboring samples.
  2. 2.
    Poor privacy-utility trade-off in existing text-to-text anonymization methods: While innovative, Feyisetan et al. (2020); Carvalho et al. (2023); Mattern et al. (2022b); Utpala et al. (2023) encounter challenges in achieving a good privacy-utility tradeoff under practical privacy budgets.
  3. 3.
    Absence of privacy budgets: The absence of detailed reporting on exact privacy budgets in (Utpala et al., 2023) hinders direct comparisons with our work.
A qualitative comparison between text-to-text privatization methods and our method is shown in Tb. 34.
Table 34: A qualitative comparison between Aug-PE and text-to-text privatization approaches.
Name Method Source of Randomness Sensitivity Control Privacy Guarantee
Madib (Feyisetan et al., 2020) TEM (Carvalho et al., 2023) Word embedding perturbation Word Embedding Noise N/A (metric distance is included in privacy definition) Word-level metric-$\epsilon d$ DP where $d$ is the distance metric
Paraphraser (Mattern et al., 2022b) DP Prompt (Utpala et al., 2023) Paraphrasing with temperature Temperature in the next token sampling stage Clipped logits of each token (Sample-level) $\epsilon$ Local DP
Aug-PE (Ours) Private evolution Histogram noise Each private sample only contributes one vote in the histogram (Dataset-level) standard $(\epsilon,\delta)$-DP
Table 35: Randomly sampled synthetic data from Madib (Feyisetan et al., 2020) (word level metric-DP $\epsilon=10$) and Aug-PE (DP $\epsilon=1$). Aug-PE with data generator GPT-3.5 yields higher quality texts.
Method Privacy Guarantee Yelp OpenReview PubMed
Madib Word level metric-DP $\epsilon=10$ i was born including raised he hardwick create during combination school , tony jones was in ’go to ’ place for the greatest pizza ever . n’t do n’t live completed whitley rich and crazy visit a preparing times a year with weeks night cut 6 civil us took to tony ’s work dinner bogota of normal end my ’local ’ strangers , tried would suggest instance visitors ? seen ’d was yet to gone …(omitted) . paper proposed bringing reinforcement buddhism based approach money automatically predictions graph augmentations for a graph neural network ( gnn ) classification problem there few authors creates taken label invariance ( data augmentations that do protect risks labels ) is part rich also felonies problem dealt gnn partner with…(omitted) mandibular overdentures many a selection treatment option for placed edentulous diabetes only long-term predictable outcomes , using suspension loading facilitated cone currently , could early well repatriation loading protocols same mandibular implant overdentures number prevalent in in literature details a systematic review ,… (omitted)
Aug-PE (GPT-3.5) DP $\epsilon=1$ The fried chicken and the collard greens were some of the best Southern fare we’ve ever had, not to mention the amazing gumbo. We highly recommend this restaurant if you’re in the area and can’t wait to try some of their other flavorful dishes. Everything filled us up and left us satisfied. This paper presents an innovative method of deep representational learning for facial expression recognition. The method is evaluated on CIFAR-10 and ImageNet datasets and it is comprehensive encompassing all facets of saliency modeling to proposed deep representational features for representing multiple saliencies. The paper is in a well-structured and the methods are clear… (omitted) In this retrospective study, we aimed to investigate the prevalence of stroke and identify the factors associated with its occurrence. Data were extracted from medical records, along with symptoms, electrocardiograms (ECGs), and syncope in a cohort of patients with a mean age of 71 years. Of the total 345 patients, 28% had cardiac abnormalities as revealed by ECGs, significantly higher than those without [p<0.001].. (omitted)
Next, we compare Aug-PE with the text-to-text privatization frameworks in detail: word-level metric-DP frameworks (Feyisetan et al., 2020; Carvalho et al., 2023) and sample-level local-DP frameworks (Mattern et al., 2022b; Utpala et al., 2023).
C.11.1 Comparison to word-level metric-DP frameworks
Madib (Feyisetan et al., 2020) and TEM (Carvalho et al., 2023) employ metric differential privacy to privatize each word independently and achieve word-level $\epsilon d$-Metric DP, where $d$ is the distance metric for neighboring words. Specifically, they perturb the embedding of each word and replace the current word with a new word whose embedding is closest to the noisy embedding. However, Aug-PE focuses on generating synthetic datasets with stronger guarantees provided by standard $(\epsilon,\delta)$-DP. Due to the fundamental differences in privacy definition: (1) metric-DP v.s. DP; (2) word-level v.s. dataset-level privacy, directly comparing our work with word-level metric-DP frameworks (Feyisetan et al., 2020; Carvalho et al., 2023) is not feasible.
To understand their privacy-utility tradeoff, we run Madib (Feyisetan et al., 2020) to generate samples under word level metric-DP with a high privacy budget $\epsilon=10$. We followed their approach of perturbing $50$-dimensional Euclidean GloVe embeddings with Laplace noise. We are unable to evaluate TEM (Carvalho et al., 2023) given that its code is not open-sourced.
Tb. 35 shows randomly sampled generated sentences from Madib and Aug-PE. Even with a high metric-DP budget ($\epsilon=10$), Madib struggles to generate meaningful sentences on Yelp, OpenReview, and PubMed datasets. In contrast, Aug-PE, with a low DP budget ($\epsilon=1$), can leverage GPT-3.5 as a data generator to produce fluent sentences across all three datasets.
C.11.2 Comparison to sample-level local-DP frameworks
Paraphraser (Mattern et al., 2022b) and DP Prompt (Utpala et al., 2023) focus on generating paraphrases for each private sample by varying the temperature during token sampling, which is regarded as a form of noise injection under the Local DP (LDP) framework. The sensitivity of each sample to the output can be constrained by clipping the logits of each generated token. While innovative, these methods’ privacy budget scales linearly with the output’s token length, presenting a challenge for generating longer sequences under a meaningful privacy budget.
It is worth noting that the mechanism for Local DP (taking a sample as input) and the mechanism for DP (taking a dataset as input) are not directly compatible. To establish a fair comparison between the Local DP in (Mattern et al., 2022b; Utpala et al., 2023) and DP employed by Aug-PE, we leveraged the conversion methodology in Feldman et al. (2022) to convert $(\epsilon_{0})$-LDP mechanism to $(\epsilon,\delta)$-DP mechanism for $\epsilon\ll\epsilon_{0}$, which requires shuffling the LDP outputs from each sample.
We use the code implementation provided by Feldman et al. (2022). [^14] https://github.com/apple/ml-shuffling-amplification Due to the constraint that $\epsilon\ll\epsilon_{0}$, [^15] https://github.com/apple/ml-shuffling-amplification/blob/993d285a546114bf8c70c33d053dca322a755707/computeamplification.py#L160 the maximal Local DP $\epsilon_{0}$ that can be used for a valid conversion on Yelp (with 1.9M private samples) is $\epsilon_{0}=8.785$, which corresponds to DP $\epsilon=1.10$.
[…]
Furthermore, our attempt to directly evaluate Paraphraser (Mattern et al., 2022b) and DP Prompt (Utpala et al., 2023) was hindered by several practical challenges.
  1. 1.
    Paraphraser: The dataset used for finetuning in Paraphraser is not publicly available, and the implementation details necessary for replicating the exact privacy guarantees are absent.
  2. 2.
    DP Prompt does not specify the exact $\epsilon_{0}$ used in the paper, focusing instead on empirical privacy attack accuracy as a proxy. The epsilon values are reported for all other baselines but not for Paraphraser and DP Prompt in Section 4.1 of Utpala et al. (2023). Additionally, as mentioned in Section 4.3 of Utpala et al. (2023), models like ChatGPT do not expose logits, so the authors do not perform a logit clipping operation in many of their experiments. This further disables the computation of an exact $\epsilon_{0}$ and renders a direct quantitative comparison between DP Prompt and Aug-PE infeasible.
C.12 Aug-PE Convergence under One Private Sample